OpenStudy (anonymous):

Someone explain or help? A circle has the following equation: x minus 1 quantity squared plus all of y minus 6 the whole squared is equal to 16 over 9 Which statement is true for the circle? Its center is at the ordered pair of 3 over 2, minus 6, and the radius is 16 over 9. Its center is at the ordered pair of 3 over 2, minus 6, and the radius is 4 over 3. Its center is at the ordered pair of minus 3 over 2, 6, and the radius is 4 over 3. Its center is at the ordered pair of minus 3 over 2, 6, and the radius is 16 over 9.

OpenStudy (anonymous):

@ nenadmatematika please do you know how to do this?

OpenStudy (anonymous):

\[(x-a)^{2}+(y-b)^{2}=r ^{2}\] this is equattion of circle with center at (a,b) and radius r

OpenStudy (campbell_st):

the equation is \[(x -1)^2 + (y - 6)^2 = \frac {16}{9}\] the general form of a circle is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre of the circle and r is the radius. find the centre and take the square root of r^2 to find the radius

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

I think..

OpenStudy (campbell_st):

reading your question the centre is (1, 6) and the radius r = 4/3

OpenStudy (campbell_st):

which doesn't match any of your listed answers

OpenStudy (anonymous):

you've lost me...so the questions is wrong?

OpenStudy (anonymous):

nooo you read it wrong lol

OpenStudy (campbell_st):

either the question is wrong or the answers are wrong...

OpenStudy (anonymous):

@campbell_st is right the answers are all wrong, or the question information is wrong

OpenStudy (anonymous):

its (x+3/2)*2 + (y+6)*2=16/9

OpenStudy (nenadmatematika):

I thought I was blind for a moment because I couldn't see the right answer...so it's possible you didn't write it correctly :D

OpenStudy (anonymous):

ahhh

OpenStudy (campbell_st):

well thats different to the question you typed.

OpenStudy (anonymous):

lol i copied and pasted it...from my worksheet.

OpenStudy (anonymous):

but anyway, no answer match it, :)

OpenStudy (anonymous):

well how would you do it...cause i still need help..

OpenStudy (campbell_st):

same problem as before no matches

OpenStudy (anonymous):

ill rewrite all the answer

OpenStudy (anonymous):

this last equation would have center at (-3/2,-6) and radius =4/3

OpenStudy (anonymous):

lol that was one of answer chooses! I chose C in the beginning and it was right lol when i was first talking to myko

OpenStudy (campbell_st):

perhaps you should read a little about the equation of a circle... it may help